OpenStudy (anonymous):

I have a quick question. I am currently studying for my math semester exam which I will take tomorrow. In the review, I have this question that I got wrong: They give me the y intercept which is 1500 and a point (2,1600) The answer is supposed to be: =63x+1500 But I don't understand how, I am pretty sure is something easy but I think I was taught this early on the semester and I have forgotten. Please help?

OpenStudy (freckles):

i assume you are looking for a line with y-int=1500 and a point (2,1600)

OpenStudy (freckles):

you can find the slope and then use point slope form to find the line

OpenStudy (freckles):

y-int=1500 or in ordered pair form (0,1500)

OpenStudy (freckles):

so find the slope of a line going through (0,1500) and (2,1600)

OpenStudy (perl):

The equation of a line has the form: y = mx + b Here m = slope, b = y intercept. you are given b = 1500 , but not the slope. so lets plug that in y = mx + 1500 . Now plug in the point (2,1600) 1600 = m(2) + 1500 solve for m

OpenStudy (freckles):

\[m=\frac{y_2-y_1}{x_2-x_1}\]

OpenStudy (freckles):

or that way

OpenStudy (perl):

solving for m : 1600 = m*2 + 1500 1600 - 1500 = 2m 100 = 2m m = 100/2

OpenStudy (anonymous):

Oh I missed that small part of (0,1500) I see

OpenStudy (perl):

the points given give us the equation y = 50x + 1500 , not y = 63x + 1500 so theres a typo in the directions

OpenStudy (anonymous):

Thank You very much for your help :) I understand now